OpenStudy (anonymous):

How do I solve x^6 when its under a square root and then under a cubed root? 3√(√x^6)

OpenStudy (igreen):

\(\sqrt[3]{\sqrt{x^6}}\) Like this?

OpenStudy (igreen):

Remember. x^6 = x * x * x * x * x * x Square root will divide the exponent by 2. 6 / 2 = ?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I just don't understand what I'm supposed to do because of the cubed root on the outside

OpenStudy (anonymous):

Would the answer just be x then?

OpenStudy (rational):

you should get |x| i think

OpenStudy (rational):

answer is just x if we assume x is positive

OpenStudy (rational):

otherwise it is |x|

OpenStudy (anonymous):

Thank you :)

OpenStudy (igreen):

Yep. Cube root would divide it by 3. \(x^6\) Square root: \(x^{6 \div 2} \rightarrow x^3\) Cube root: \(x^{3 \div 3} \rightarrow x\)

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'. @nicf